# Bayes optimal learning of attention-indexed models

### v2

@@ C.3 Output Channel Computation
$\widehat{q}_{\ell k}=\frac{4\alpha}{q_{\ell k}}\sum_{a\leq b}\int\!\mathcal{D}\eta\,dy\;\frac{\bigl(\partial_{\eta_{ab}^{(\ell)}}\mathcal{Z}_{\mathrm{out}}\bigr)\,\bigl(\partial_{\eta_{ab}^{(k)}}\mathcal{Z}_{\mathrm{out}}\bigr)}{\mathcal{Z}_{\mathrm{out}}}.$ (126)
Recalling the definition of the output denoising function, we can express the output channel state equation of our model as:
@@
$\begin{split}\widehat{q}_{\ell k}&=4\alpha\;\mathbb{E}_{\eta,y}\!\left[\sum_{a\leq b}(g_{\rm out}(y,\omega,V))_{ab}^{(\ell)}(g_{\rm out}(y,\omega,V))_{ab}^{(k)}\right],\\ &\quad\text{for}\quad\ell=1,\dots,L\quad a\leq b=1,\dots,T\end{split}$ (127)
@@
$P_{\rm out}\!\Bigl(y\,\Bigm|\,\{h_{ab}^{(\ell)}\}_{\ell}\Bigr)=\delta(\{y_{ab}-g(\{h^{(\ell)}_{ab}\}_{\forall\ell})_{ab}\}_{\forall ab})$ (130)
@@
$P_{\rm out}\!\Bigl(y\,\Bigm|\,\{h_{ab}^{(\ell)}\}_{\ell}\Bigr)=\delta(\{y_{ab}-B^{L}_{c}(\{h^{(\ell)}_{ab}\}_{\forall\ell})_{ab}\}_{\forall ab})$ (131)
@@ C.4 The fixed point of AMP is described by the state equations
We now claim, in analogy with [zdeborova2016statistical, troiani2025fundamentallimitslearningsequence, maillard_bayes-optimal_2024, erba2024bilinear] that for every sample $\mu$ and every couple of tokens $a\leq b$ the variables $\omega_{\mu,\,ab}^{t}$ at each time converge to independent centered Gaussian variables with the following covariances
@@
By Nishimori’s identities [Nishimori_1980] we can assume $m^{t}=q^{t}$. The first equation of (21) is now immediately recovered (modulo the substitution $V\to 2(Q-q^{t})$ which will come after)
@@
Again as in [zdeborova2016statistical, troiani2025fundamentallimitslearningsequence, maillard_bayes-optimal_2024, erba2024bilinear] we will have that in distribution
@@ D.1 Prior channel state equation
where $Q=1+\rho$.Then, one has the relation from [maillard2022perturbative]:
@@ D.2.1 Small width limit
We follow [maillard_bayes-optimal_2024, Section E.1.1]. Call $t=\rho/{\hat{q}}$, and ${\bar{\alpha}}=\alpha/\rho$. Call $\nu$ the distribution of $\sqrt{\rho}(S_{*}+\frac{1}{\sqrt{{\hat{q}}}}Z)=\sqrt{\rho}S_{*}+\sqrt{t}Z$, i.e.
@@
Notice that this is precisely the $\nu$ defined in [maillard_bayes-optimal_2024, Eq. 56]. Then we have
@@
where we used [maillard_bayes-optimal_2024, Eq. 57 and following] to take the limit of small $\kappa$ at leading order. Thus, the equations can be recast to
@@ D.2.2 Large width limit
where we used [maillard_bayes-optimal_2024, Section E.2].

### v3

@@ C.3 Output Channel Computation
$\frac{\partial I_{\rm out}(q)}{\partial q_{\ell k}}=\int\Bigl[\prod_{a\leq b}d^{L}\omega_{ab}\,\mathcal{N}(\omega_{ab};0,2q)\Bigr]\;\left(\sum_{a\leq b}\frac{\partial^{2}F(\omega,q)}{\partial\omega^{(\ell)}_{ab}\,\partial\omega^{(k)}_{ab}}\;+\;\frac{\partial F(\omega,q)}{\partial q_{\ell k}}\Big|_{\omega}\right),$ (127)
where $\partial F/\partial q|_{\omega}$ is the explicit derivative through V=2(Qq)V=2(Q-q) at fixed $\omega$.
Since $V^{(rs)}=2(Q_{rs}-q_{rs})$, we have $\partial V^{(rs)}/\partial q_{\ell k}=-2\delta_{r\ell}\delta_{sk}$, hence
$\frac{\partial F(\omega,q)}{\partial q_{\ell k}}\Big|_{\omega}=-2\,\frac{\partial F(\omega,q)}{\partial V^{(\ell k)}}.$ (128)
Plugging (128) into (127) gives
$\frac{\partial I_{\rm out}(q)}{\partial q_{\ell k}}=\int\Bigl[\prod_{a\leq b}d^{L}\omega_{ab}\,\mathcal{N}(\omega_{ab};0,2q)\Bigr]\;\left(\sum_{a\leq b}\frac{\partial^{2}F}{\partial\omega^{(\ell)}_{ab}\,\partial\omega^{(k)}_{ab}}\;-\;2\,\frac{\partial F}{\partial V^{(\ell k)}}\right).$ (129)
Now fix $\omega,V$ and define for brevity
$Z(y):=\mathcal{Z}_{\rm out}(y,\omega,V),\qquad g_{ab}^{(\ell)}(y):=(g_{\rm out}(y,\omega,V))_{ab}^{(\ell)}.$ (130)
Then
$F(\omega,q)=\int dy\;Z(y)\,\ln Z(y).$ (131)
We compute the two derivatives in (129) exactly. Since $Z$ is an integral of a Gaussian density $\mathcal{N}(h_{ab};\omega_{ab},V)$, then by definition:
$\frac{\partial Z(y)}{\partial\omega^{(\ell)}_{ab}}=Z(y)\,g^{(\ell)}_{ab}(y),\qquad\frac{\partial^{2}Z(y)}{\partial\omega^{(\ell)}_{ab}\partial\omega^{(k)}_{ab}}=Z(y)\,g^{(\ell)}_{ab}(y)\,g^{(k)}_{ab}(y)+Z(y)\,\frac{\partial g^{(\ell)}_{ab}(y)}{\partial\omega^{(k)}_{ab}}$. (132)
Therefore,
$\frac{\partial F}{\partial\omega^{(k)}_{ab}}=\int dy\;(1+\ln Z(y))\,\frac{\partial Z(y)}{\partial\omega^{(k)}_{ab}}=\int dy\;(1+\ln Z(y))\,Z(y)\,g^{(k)}_{ab}(y),$ (133)
$\frac{\partial^{2}F}{\partial\omega^{(\ell)}_{ab}\partial\omega^{(k)}_{ab}}=\int dy\;\Bigl[Z(y)\,g^{(\ell)}_{ab}(y)\,g^{(k)}_{ab}(y)+(1+\ln Z(y))\Bigl(Z(y)\,g^{(\ell)}_{ab}(y)\,g^{(k)}_{ab}(y)+Z(y)\,\partial_{\omega^{(k)}_{ab}}g^{(\ell)}_{ab}(y)\Bigr)\Bigr].$
For a Gaussian integral, the derivative of $\ln Z$ with respect to $V$ satisfies the identity
$\frac{\partial\ln Z(y)}{\partial V^{(\ell k)}}=\frac{1}{2}\left(g^{(\ell)}_{ab}(y)\,g^{(k)}_{ab}(y)+\frac{\partial g^{(\ell)}_{ab}(y)}{\partial\omega^{(k)}_{ab}}\right)$, (135)
Using $\partial_{V^{(\ell k)}}Z=Z\,\partial_{V^{(\ell k)}}\ln Z$ and (135), we obtain
$\frac{\partial F}{\partial V^{(\ell k)}}=\int dy\;(1+\ln Z(y))\,\frac{\partial Z(y)}{\partial V^{(\ell k)}}=\frac{1}{2}\int dy\;(1+\ln Z(y))\,Z(y)\,\left(g^{(\ell)}_{ab}(y)\,g^{(k)}_{ab}(y)+\partial_{\omega^{(k)}_{ab}}g^{(\ell)}_{ab}(y)\right).$
Subtracting $2\,\partial F/\partial V^{(\ell k)}$ from $\partial^{2}F/(\partial\omega^{(\ell)}_{ab}\partial\omega^{(k)}_{ab})$ using (134) and (136), all terms proportional to $(1+\ln Z)Z(\cdots)$ cancel exactly, yielding
$\frac{\partial^{2}F}{\partial\omega^{(\ell)}_{ab}\partial\omega^{(k)}_{ab}}-2\,\frac{\partial F}{\partial V^{(\ell k)}}=\int dy\;Z(y)\,g^{(\ell)}_{ab}(y)\,g^{(k)}_{ab}(y).$ (137)
Plugging (137) into (129) and summing over $a\leq b$ gives
$\frac{\partial I_{\rm out}(q)}{\partial q_{\ell k}}=\int\Bigl[\prod_{a\leq b}d^{L}\omega_{ab}\,\mathcal{N}(\omega_{ab};0,2q)\Bigr]\;\sum_{a\leq b}\int dy\;\mathcal{Z}_{\rm out}(y,\omega,V)\,(g_{\rm out}(y,\omega,V))_{ab}^{(\ell)}(g_{\rm out}(y,\omega,V))_{ab}^{(k)}.$
Therefore the output-channel state equation is
$\widehat{q}_{\ell k}=4\alpha\;\mathbb{E}_{\omega,y}\!\left[\sum_{a\leq b}(g_{\rm out}(y,\omega,V))_{ab}^{(\ell)}\,(g_{\rm out}(y,\omega,V))_{ab}^{(k)}\right],$ (139)
which is equivalent to
@@
$\begin{split}\widehat{q}_{\ell k}&=4\alpha\;\mathbb{E}_{\eta,y}\!\left[\sum_{a\leq b}(g_{\rm out}(y,\omega,V))_{ab}^{(\ell)}(g_{\rm out}(y,\omega,V))_{ab}^{(k)}\right],\\ &\quad\text{for}\quad\ell=1,\dots,L\quad a\leq b=1,\dots,T.\end{split}$ (140)
@@
$P_{\rm out}\!\Bigl(y\,\Bigm|\,\{h_{ab}^{(\ell)}\}_{\ell}\Bigr)=\delta(\{y_{ab}-g(\{h^{(\ell)}_{ab}\}_{\forall\ell})_{ab}\}_{\forall ab}),$ (143)
@@
$P_{\rm out}\!\Bigl(y\,\Bigm|\,\{h_{ab}^{(\ell)}\}_{\ell}\Bigr)=\delta(\{y_{ab}-B^{L}_{c}(\{h^{(\ell)}_{ab}\}_{\forall\ell})_{ab}\}_{\forall ab}).$ (144)
@@ C.4 The fixed point of AMP is described by the state equations
We now claim, in analogy with [51, 43, 32, 20] that for every sample $\mu$ and every couple of tokens $a\leq b$ the variables $\omega_{\mu,\,ab}^{t}$ at each time converge to independent centered Gaussian variables with the following covariances
@@
By Nishimori’s identities [37] we can assume $m^{t}=q^{t}$. The first equation of (21) is now immediately recovered (modulo the substitution $V\to 2(Q-q^{t})$ which will come after)
@@
Again as in [51, 43, 32, 20] we will have that in distribution
@@ D.1 Prior channel state equation
where $Q=1+\rho$.Then, one has the relation from [30]:
@@ D.2.1 Small width limit
We follow [32, Section E.1.1]. Call $t=\rho/{\hat{q}}$, and ${\bar{\alpha}}=\alpha/\rho$. Call $\nu$ the distribution of $\sqrt{\rho}(S_{*}+\frac{1}{\sqrt{{\hat{q}}}}Z)=\sqrt{\rho}S_{*}+\sqrt{t}Z$, i.e.
@@
Notice that this is precisely the $\nu$ defined in [32, Eq. 56]. Then we have
@@
where we used [32, Eq. 57 and following] to take the limit of small $\kappa$ at leading order. Thus, the equations can be recast to
@@ D.2.2 Large width limit
where we used [32, Section E.2].